Building a flower bed for the wife and it's been years since I've done calculations.
Question: How many cubic feet in a round flower bed 10 feet in diameter, 6 inches deep?

Answer by jim_thompson5910(35256)

Volume = pi*r^2*h where r is the radius and h is the height

V = pi*r^2*h

V = (3.14159)*(5^2)*(0.5) *** See Note Below ***

V = (3.14159)*(25)*(0.5)

V = 39.269875

So the volume of the flower bed is roughly 39.269875 cubic feet.


Note: Make sure everything is in the same units. So we want cubic feet, which means it makes sense to get everything into feet. So 6 inches = 0.5 feet

Also, if you have a diameter of 10 feet, then the radius is 10/2 = 5 feet
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------